OpenStudy (anonymous):

Use linear approximation (or differentials) to estimate the given number: 3 square root 1001

OpenStudy (anonymous):

\[\sqrt[3]{1001}\]

OpenStudy (dumbcow):

for function f(x) the linear approximation at point "a" is: \[f(x) \approx f'(a)(x-a) + f(a)\] for this problem f(x) = x^1/3 let a = 1000